OpenStudy (anonymous):

|2-x|/2-x = 0 How do I get the 2-x thats under the fraction over to the other side

OpenStudy (anonymous):

just multiply it to both sides.

OpenStudy (anonymous):

So would i multiply it by 2+x or 2-x?

OpenStudy (anonymous):

well you want to cancel the 2-x so you multiply that to both sides.

OpenStudy (anonymous):

because that will get you 2-x/2-x on the left side and thats just equal to 1

OpenStudy (anonymous):

Then I would get |2-x|(2-x) = 0 wouldnt I?

OpenStudy (anonymous):

You don't. What's under the fraction cannot equal zero (unless there's a limit around there). That means x can take any value from your domain ( I'll assume that would be R) except 2 (because division by zero has been outlawed centuries ago). Because what's under the fraction can't equal zero, it means that what's above the fraction has to. However, in order for |x-2| to equal zero, x has to equal 2. But that goes against the previous paragraph. Which makes the exercise redundant.

OpenStudy (anonymous):

He doesn't want what's under the fraction to be zero.

OpenStudy (anonymous):

*The point of my previous post was that the exercise itself is redundant. Any attempt to solve it will fail because it's non-sense.

OpenStudy (anonymous):

|2-x|/2-x = 0 ima multiply both sides by 2-x (2-x)|2-x|/(2-x) = 0*(2-x)

OpenStudy (anonymous):

in the left side you have 2-x/2-x what happens when you have the same thing divided by itself for example 2/2 or 4/4 they equal 1

OpenStudy (anonymous):

Ohh okay, i thought there was more steps than that, I was just overthinking it. Thank you!

OpenStudy (anonymous):

yeah in the end you are left with |2-x|=0

OpenStudy (anonymous):

which pretty much yields you one solution for x

OpenStudy (anonymous):

*facepalm* Which will get you to x=2 which will go against your inition |2-x| / (2-x) because you're dividing by zero. Which is non-sense, wrong and bad.

OpenStudy (anonymous):

So i typed it into my hw and it tells me that the x-intercept does not exist

terenzreignz (terenzreignz):

lolz for any value that a sane person would substitute for x in the expression \[\Large \frac{|2-x|}{2-x}\] the result would always be either 1 or -1, never anything else. Imagine y = 2-x, so that \[\Large \frac{|y|}y\] and it'd probably make more sense that y cannot be 0, and neither can 2-x

OpenStudy (anonymous):

So are you saying |2-x|=0 at the end is not the answer

terenzreignz (terenzreignz):

While it's tempting to just multiply both sides by 2-x at the beginning, don't forget that in every instance that you do that, say in algebra, or the like, you are assuming from the start that what you multiplied to both sides was NOT zero to begin with. Because if you multiply zero to both sides, the resulting would ALWAYS be equal. So, implicitly, what was done was this: \[\Large \frac{|2-x|}{2-x}\times (2-x) = 0\times (2-x) \ \ assuming \ 2 -x \ne 0\] \[\Large |2-x| = 0 \ \ \ \ a \ contradiction\] Never mind these details for now, though. Just accept that there must NEVER be a zero in the denominator.

OpenStudy (anonymous):

I understand that 2-x cannot equal zero and maybe I should have explained the question more. They give me |2-x| / 2-x and want me to graph it. So I need to do the piecewise function for {-x if x<0 , x>0} and thats why I was trying to get the problem into |absolute value|=c

terenzreignz (terenzreignz):

Then I've already pretty much made it a dead giveaway :P There are three cases then, one of which was discussed here in heated detail, and it's that if x = 2, the function does not exist, yes? So, let's put that it. \[\Large f(x) = \begin{cases}???&x < 2\\ \emptyset & x = 2\\??? &x>2\end{cases}\] So... I want you to do something for me (actually yourself, LOL) Try plugging in random numbers that are less than 2, see what you get.

OpenStudy (anonymous):

Where do I plug them in at? When the equation is still |2-x|/2-x ?

terenzreignz (terenzreignz):

Yup. Plug in certain number that are less than 2, say, 1, 0, -1, your choice, it can be as mind bogglingly 'complex' as \(\large -\pi\) for all I care XD

OpenStudy (anonymous):

I get 1 each time

terenzreignz (terenzreignz):

No you don't?

OpenStudy (anonymous):

|2-1|/2-1 = 1/1 =1 |2-0|/2-0 = 2/2 = 1

terenzreignz (terenzreignz):

Oops. Yeah, you're right :3 okay, try plugging in numbers greater than 2 now

OpenStudy (anonymous):

I get 1 again

terenzreignz (terenzreignz):

Are you sure? :P

OpenStudy (anonymous):

|2-3| / 2-3 = -1/-1 = 1

OpenStudy (anonymous):

oh wait its a positive on top isnt it

terenzreignz (terenzreignz):

There we go XD

terenzreignz (terenzreignz):

so in fact, for all x greater than 2, the function is simply equal to...?

OpenStudy (anonymous):

-1?

terenzreignz (terenzreignz):

That's right. So, in fact, for all x less than 2, the function is just equal to 1 for x = 2, the function has no value, so leave that as a hole of sorts for x > 2, the function is equal to -1 Should be simple to graph now ^_^ \[\Large f(x) = \begin{cases}1&x < 2\\ \emptyset & x = 2\\-1&x>2\end{cases}\]

OpenStudy (anonymous):

Thank you so much!! :)

terenzreignz (terenzreignz):

No worries.